# How do Probabilistic Graphical Models and Graph Neural Networks Look at Network Data?

### v2

@@ 3 Background
We imagine a scenario when a practitioner is willing to analyze a network problem and its corresponding dataset, and is agnostic on the model family GNN vs. PGM. Hence, we consider here off-the-shelf implementations of various main models in each of these families, and refrain from using models specifically designed for one main task, e.g. specifically designed for handling heterophilic or disassortative graphs. This also mimics the realistic scenario where the practitioner does not know in advance if their dataset exhibit a specific structure.
@@ 3.2 GNN architectures
The choices of the functions $\phi^{(\ell)}$, $g^{(\ell)}$ differ based on the model. Contrarily to PGMs, they usually entail non-linearities. In our experiments, we employ three distinct architectures: Graph Attention Networks (GAT) [27], which use an attention mechanism inside the function $g$; Graph Autoencoders (GAE) [28], and Graph Variational Autoencoders (VGAE) [28], which use an autoencoder framework to learn embeddings, with GAE being a non-probabilistic variant of VGAE. The cost function in VGAE contains a KL divergence term, together with the binary cross-entropy.
@@ 4.1 Datasets
We generate synthetic networks with community structure. We assign $K$-dimensional membership vectors to nodes, assuming an underlying community structure consisting of $K$ overlapping groups. We then draw a $K\times K$ affinity matrix controlling the density of edges between pairs of nodes. This follows the generative models behind the PGMs considered here.
Specifically, we generate directed networks of $N\in\left\{100,1000\right\}$ number of nodes; $K=5$ communities and average degree $\langle k\rangle\in\left\{5,20\right\}$. To control the heterophily of the network, we manipulate the density of edges between nodes in different communities. Heterophilic have larger off-diagonal elements of the affinity matrix. For each combination of size and average degree, we generate one strong homophilic and one strong heterophilic network. We sample $10$ networks for each parameter configuration.
@@
Table 1: Dataset statistics for the 10 real datasets analysed. $N$ is the number of nodes; $\langle k\rangle$ is the average degree of a node; $h$ is the edge homophily ratio (i.e., the fraction of edges connecting nodes with the same class label, computed only for datasets with node labels); $F$ is the dimension of the node feature (when available).
@@ 5 The role of input features
In case 1)1)1 ), we assign to each node $i$ the corresponding row in the adjacency matrix $A_{i}=[A_{i1},\dots,A_{iN}]$. This information is readily available as it is a quantity simply derived from the input matrix $A$. We stress that, when splitting the graph in training, validation and test set in our cross-validation link prediction routine, we take into account the masking of the edges, and we feed the node with the masked adjacency matrix row. Indeed, the training set feature only contains the links between nodes in the training set graph, while for validation and test we progressively include the validation and test set links.
In case 2)2)2 ), the features are directly available only for 7 out of 10 real datasets. For the real datasets that do not come equipped with engineered features, we assess their performance by simply feeding the structure-based feature.
@@
Figure 1: Structure-based vs. attribute-based features. AUC scores of link prediction in real dataset when giving in input different types of features. For the PGMs, only MTCOV uses input features. Bars and errors are mean and standard deviation over 5 cross-validation folds.
@@
Figure 2: High vs. low dimensional features. AUC scores of link prediction in real and synthetic dataset when giving in input either the original feature of dimension $F$ (Full) or the one clustered with K-means (Clustered), i.e. $F=1$. For the PGMs, the two bars correspond to the same AUC value, given that MTCOV takes one-dimensional (Clustered) features by design and the other two PGMs no feature at all. Bars and errors are mean and standard deviation over 5 cross-validation folds.
@@ 6 Robustness to noisy features
We find that PGMs are more robust when increasing the percentage of nodes whose features are randomized, highlighting the stronger reliance of GNNs on input feature information, see Figure 3. In particular, GAT appears to be more sensitive to noise compared to autoencoder-based architectures (VGAE, GAE). This enhanced sensitivity may be related to the fact that, when features contain noise, the attention scores can be misleading, and we notice this is especially pronounced when the input features have lower dimensionality, as in PPI ($F=50$) and Peptides ($F=9$). While VGAE and GAE generally exhibit greater resilience to noise, they become more sensitive when the dataset is heterophilic, as in Wisconsin and Chameleon.
Figure 3: AUC score variation in real dataset when increasing the percentage of nodes $\rho$ whose features are randomized. The AUC score variation is calculated as the difference between the $AUC_{\rho}$ and the original $AUC_{0}$, corresponding to the case where $0\%$ of nodes have randomized features. Markers and errors are mean and standard deviation over 5 cross-validation folds.
@@ 7 Resilience to increasing heterophily
Having studied in detail the role of input features in learning missing links, we now focus on comparing the two frameworks for different network clustering structure. Specifically, the third experiment explores the impact of increasing the heterophily (or disassortativeness) of the graph. This is the property of observing nodes in different clusters more likely to be connected, which contrasts with homophilic (or assortative) networks, where we expect nodes in the same cluster to have higher chances of being connected. In PGMs, learning in disassortative networks has often been shown to be a more challenging task than in assortative networks [21, 45, 46]. Similar declines in performance in heterophilc networks are also now increasingly investigated in GNNs [10, 11, 12, 13, 14, 15, 16].
@@
In the synthetic case, we can directly manipulate the level of heterophily by varying the density of edges between nodes in different clusters in the affinity matrix [21]. In real data, when extra node labels are available, and assumed to be informative, we measure the homophily edge ratio $h$ as the fraction of edges in a graph which connect nodes that have the same class label, as done e.g. in [13]. A decreasing value of $h$ indicates increased heterophily. Four of the real datasets have $h<0.4$.
We find that PGMs significantly outperform GNNs when increasing the heterophily of the graph in all datasets (both synthetic and real), when given the full input feature, see Figure 4. In particular, autoencoder-based architectures (VGAE, GAE) suffer a significant decrease in performance in two datasets (Wisconsin and Texas). However, we also find that by giving in input to the GNN the row of the adjacency matrix (structure-based feature) instead of the attribute-based feature, the situation changes favorably. In this case, GNNs show similar performance to PGMs in two datasets (Email and Chameleon). This result highlights that while GNNs are expected to be generally less robust in heteropilic graphs, performance could be improved by a better choice of input features, rather than acting on a change in architecture design as proposed in [11, 47, 48]. A similar conclusion was also observed in node classification tasks [49], where GNN architectures were shown to significantly improve their performance under good conditions of heterophily, specifically when nodes sharing the same label also exhibit similar embeddings and neighbor distributions. Moreover, the phenomenon of oversmoothing [50, 51, 52, 53], which is commonly associated with poor performance on heterophilic graphs [11, 49, 54], has been shown to be avoidable without altering the GNN architecture, simply by increasing the variance of weight initialization [55].
@@
Figure 4: AUC scores of link prediction in real and synthetic heterophilic dataset when giving in input the full feature. For the real dataset, we compare the AUC when feeding the GNN with structure-based or attribute-based features. For the PGMs, only MTCOV uses input features. Bars and errors are mean and standard deviation over 5 cross-validation folds.
@@ 8.1 Interpretability
To create a parallelism with the interpretation of communities in PGMs and the output GNNs’ embeddings, we visualize the memberships and embeddings in a two-dimensional plan by applying a dimensionality reduction to these vectors using t-SNE [63]. We use a synthetic network as example, as we have access to ground truth (GT) communities. We first notice how nodes that have a more pronounced mixed-membership are placed on the border of their main community, closer to the other communities they also participate in, see Figure 5. This is somehow expected when plotting the GT $u_{i}$, given their shape. However, it is not trivial that the models are able to maintain this positional property in the inferred parameters. We find that both MULTITENSOR and GAT are able to place more mixed nodes closer to the border, as observed for the GT. For instance, nodes that belong to a large extent to two communities, are placed in between the two main clusters. However, while mixed-membership PGMs allow for a direct estimation of the mixed-membership, as this is their natural output, it is not clear how to extract this information automatically and in a principled way (beside visual inspection) in GNNs from their output embeddings, as they lack the innate ability to represent these memberships explicitly. To enhance the interpretability of GNNs, one promising direction is to develop methods that enable the extraction of interpretable mixed-membership information from their embeddings. Achieving this, however, would require architectural modifications to GNNs, as demonstrated in [61].
@@ 8.2 Computational Complexity
Table 2: Computational complexities for the models under analysis. $N$ the number of nodes, $E$ the total number of edges, $K$ the number of communities, $Z$ the number of attribute categories, $F$ the input feature dimension, $F^{\prime}$ the hidden layer dimension, and $L$ the number of GNN layers.
@@
9 Related Work
Our aim is to directly compare Probabilistic Graphical Models (PGMs) and Graph Neural Networks (GNNs) as distinct frameworks for handling network data. This comparison is lacking in the literature, as these two types of approaches usually concern different scientific communities, who adopt distinct languages and distinct model validation criteria, making comparisons hard.
While several studies have explored hybrid approaches combining the strengths of PGMs and GNNs [61, 64, 65, 66], such as leveraging PGMs to enhance the explainability of GNNs [65] or employing GNNs to improve the scalability and efficiency of PGMs [66], we think that a direct comparison between these two frameworks would provide deeper insights into their respective advantages and limitations.
In particular, we address the growing recognition in the literature that GNNs often struggle when provided with insufficiently informative input features and when the graph structure is heterophilic, as they cannot fully leverage graph topology alone to achieve robust performance and strongly rely on input features [67, 68]. Moreover, node clustering is often tackled in an unsupervised regime, where node labels are not available as input. In this setting, GNNs have exhibited weaker performance compared to supervised or semi-supervised approaches [58].
Even in a semi-supervised setting, GNNs have been shown to be outperformed by PGMs on a node classification task. Specifically, results indicate a significant performance gap between the Contextual Stochastic Block Model (CSBM), introduced for unsupervised community detection on attributed graphs, and state-of-the-art GNN architectures [69].
@@ 10 Conclusions
Notably, previous studies have shown that linear models can outperform GNNs in link prediction tasks [70, 71]. Given these limitations, GNNs should ideally be designed to learn automatically whether to discard irrelevant features or to select what features to use, potentially incorporating structural information such as the adjacency matrix of the graph. In the absence of this additional learning step, care should be taken when choosing what feature to give in input.
@@ Appendix A Additional plots
In Figure 7, we compare the performance of PGM and GNN models when using structure-based input features across both synthetic and real datasets. The performance of GNNs is generally comparable to that of PGMs in this setting, as shown before in Figure 1.
@@
Figure 8 further illustrates the performance of GNNs when fed either the full input feature or the clustered one, across both real and synthetic datasets. As shown before in Figure 2, we observe a significant drop in GNN performance when the input feature is clustered, and this effect is consistent across all datasets.
Figure 8: High vs. low dimensional features. AUC scores of link prediction in real and synthetic datasets when giving in input either the original feature of dimension $F$ (Full) or the one clustered with K-means (Clustered), i.e. $F=1$. For the PGMs, the two bars correspond to the same AUC value, given that MTCOV takes one-dimensional (Clustered) features by design and the other two PGMs no feature at all. Bars and errors are mean and standard deviation over 5 cross-validation folds.

### v3

@@ 1 Introduction
2 Related Work
Our aim is to directly compare Probabilistic Graphical Models (PGMs) and Graph Neural Networks (GNNs) as distinct frameworks for handling network data. This comparison is lacking in the literature, as these two types of approaches usually concern different scientific communities, who adopt distinct languages and distinct model validation criteria, making comparisons hard [17].
While several studies have explored hybrid approaches combining the strengths of PGMs and GNNs [18, 19, 20, 21, 22], such as leveraging PGMs to enhance the explainability of GNNs [20] or employing GNNs to improve the scalability and efficiency of PGMs [21], we think that a direct comparison between these two frameworks would provide deeper insights into their respective advantages and limitations.
In particular, we address the growing recognition in the literature that GNNs often struggle when provided with insufficiently informative input features and when the graph structure is heterophilic, as they cannot fully leverage graph topology alone to achieve robust performance and strongly rely on input features [23, 24]. Moreover, node clustering is often tackled in an unsupervised regime, where node labels are not available as input. In this setting, GNNs have exhibited weaker performance compared to supervised or semi-supervised approaches [25]. Notably, even in a semi-supervised setting, GNNs have been outperformed by the Contextual Stochastic Block Model (CSBM) [26] on a node classification task.
@@ 4 Background
We imagine a scenario when a practitioner is willing to analyze a network problem and its corresponding dataset, and is agnostic on the model family GNN vs. PGM. Hence, we consider here off-the-shelf implementations of various main models in each of these families, and refrain from using models specifically designed for one main task, e.g. specifically designed for handling heterophilic or disassortative graphs. This also mimics the realistic scenario where the practitioner does not know in advance if their dataset exhibits a specific structure.
@@ 4.2 GNN architectures
The choices of the functions $\phi^{(\ell)}$, $g^{(\ell)}$ differ based on the model. Contrarily to PGMs, they usually entail non-linearities. In our experiments, we employ four distinct architectures: Graph Attention Networks (GAT) [37], which use an attention mechanism inside the function $g$; Graph Autoencoders (GAE) [38] and Graph Variational Autoencoders (VGAE) [38], which use an autoencoder framework to learn embeddings, with GAE being a non-probabilistic variant of VGAE; H2GCN [39], a GNN architecture specifically designed to handle heterophily. H2GCN accounts for the fact that in heterophilic networks neighboring nodes may have dissimilar features and labels, unlike standard GNNs that assume instead similarity. It does so by specific design choices in the neural network architecture, such as separate $\phi_{i}^{(\ell)}$ and $\phi_{\partial i}^{(\ell)}$ for self and neighbor embeddings, skip connections and embeddings from second-order neighbors. These have been shown helping to mitigate the oversmoothing effect [40, 41, 42, 43], which is known to worsen GNN performance in heterophilic settings.
@@ 5.1 Datasets
We generate synthetic networks with community structure. We assign $K$-dimensional membership vectors to nodes, assuming an underlying community structure consisting of $K$ overlapping groups. We then draw a $K\times K$ affinity matrix controlling the density of edges between pairs of nodes. This follows the generative models behind the PGMs considered here. Specifically, we generate directed networks of $N\in\left\{100,1000\right\}$ number of nodes; $K=5$ communities and average degree $\langle k\rangle\in\left\{5,20\right\}$. To control the heterophily of the network, we manipulate the density of edges between nodes in different communities. Heterophilic have larger off-diagonal elements of the affinity matrix. For each combination of size and average degree, we generate one homophilic and one heterophilic network. We sample $10$ networks for each parameter configuration. More details are provided in Appendix B.
@@
Table 1: Dataset statistics for the 10 real datasets analysed. NN is the number of nodes; $\langle k\rangle$ is the average degree of a node; $h$ is the edge homophily ratio (i.e., the fraction of edges connecting nodes with the same class label, computed only for datasets with node labels); $F$ is the dimension of the node feature (when available).
@@ 6 The role of input features
In case $1)$, we assign to each node $i$ the corresponding row in the adjacency matrix $A_{i}=[A_{i1},\dots,A_{iN}]$. This information is readily available as it is a quantity simply derived from the input matrix $A$. We stress that, when splitting the graph in training, validation and test set in our cross-validation link prediction routine, we take into account the masking of the edges, and we feed the node with the masked adjacency matrix row. Indeed, the training set feature only contains the links between nodes in the training set graph, while for validation and test we progressively include the validation and test set links.
In case $2)$, the features are directly available only for 7 out of 10 real datasets. For the real datasets that do not come equipped with engineered features, we assess their performance by simply feeding the structure-based feature.
@@
Figure 1: Structure-based vs. attribute-based features. AUC scores of link prediction in real datasets when giving in input different types of features. For the PGMs, only MTCOV uses input features. Bars and errors are mean and standard deviation over 5 cross-validation folds.
@@
Figure 2: High vs. low dimensional features. AUC scores of link prediction in real and synthetic datasets when giving in input either the original feature of dimension $F$ (Full) or the one clustered with K-means (Clustered), i.e. $F=1$. For the PGMs, MTCOV takes one-dimensional (Clustered) features by design and the other two PGMs no feature at all. Bars and errors are mean and standard deviation over 5 cross-validation folds.
@@ 7 Robustness to noisy features
We find that PGMs are more robust when increasing the percentage of nodes whose features are randomized, highlighting the stronger reliance of GNNs on input feature information, see Figure 3. In particular, GAT and H2GCN appear to be overall more sensitive to noise compared to autoencoder-based architectures (VGAE, GAE). The GAT enhanced sensitivity may be related to the fact that, when features contain noise, the attention scores can be misleading, and we notice this is especially pronounced when the input features have lower dimensionality, as in PPI ($F=50$) and Peptides ($F=9$). While VGAE and GAE generally exhibit greater resilience to noise, they become more sensitive when the dataset is heterophilic, as in Wisconsin and Chameleon.
Figure 3: AUC score variation in real datasets when increasing the percentage of nodes $\rho$ whose attribute-based features are randomized. The AUC score variation is calculated as the difference between the $AUC_{\rho}$ and the original $AUC_{0}$, corresponding to the case where $0\%$ of nodes have randomized features. Markers and errors are mean and standard deviation over 5 cross-validation folds.
@@ 8 Resilience to increasing heterophily
Having studied in detail the role of input features in learning missing links, we now focus on comparing the two frameworks for different network clustering structure. Specifically, the third experiment explores the impact of increasing the heterophily (or disassortativeness) of the graph. This is the property of observing nodes in different clusters more likely to be connected, which contrasts with homophilic (or assortative) networks, where we expect nodes in the same cluster to have higher chances of being connected. In PGMs, learning in disassortative networks has often been shown to be a more challenging task than in assortative networks [31, 60, 61, 62]. Similar declines in performance in heterophilc networks are also now increasingly investigated in GNNs [10, 11, 12, 13, 14, 15, 16].
@@
In the synthetic case, we can directly manipulate the level of heterophily by varying the density of edges between nodes in different clusters in the affinity matrix [31]. In real data, when extra node labels are available, and assumed to be informative, we measure the homophily edge ratio $h$ as the fraction of edges in a graph which connect nodes that have the same class label, as done e.g. in [13]. A decreasing value of $h$ indicates increased heterophily. Four of the real datasets have $h<0.4$. In the synthetic case we compute $h$ by assigning each node the label corresponding to the argmax of its membership vector.
@@
Figure 4: AUC scores of link prediction in real and synthetic heterophilic datasets when using the full feature. For the real dataset, we compare the AUC when feeding the GNN with structure-based or attribute-based features. For the PGMs, only MTCOV uses input features. Bars and errors are mean and standard deviation over 5 cross-validation folds.
We find that PGMs generally outperform GNNs when increasing the heterophily of the graph (both synthetic and real), when given the full input feature, see Figure 4. In particular, autoencoder-based architectures (VGAE, GAE) suffer a significant decrease in performance in two datasets (Wisconsin and Texas). However, we also find that by giving in input to the GNN the row of the adjacency matrix (structure-based feature) instead of the attribute-based feature, the situation changes favorably. In this case, GNNs show similar performance to PGMs in two datasets (Email and Chameleon). This applies to all GNN models, including H2GCN, which is explicitly designed to handle heterophily. Nonetheless, even under these favorable conditions, H2GCN’s performance remains comparable to or lower than that of PGMs, suggesting that architectural adaptations alone may not fully address the challenges posed by heterophilic graphs. This result highlights that while GNNs are expected to be generally less robust in heteropilic graphs, performance could be improved by a better choice of input features, rather than acting on a change in architecture design as proposed in [11, 39, 63]. A similar conclusion was also observed in node classification tasks [64], where GNN architectures were shown to significantly improve their performance under good conditions of heterophily, specifically when nodes sharing the same label also exhibit similar embeddings and neighbor distributions. Moreover, the phenomenon of oversmoothing [40, 41, 42, 43], which is commonly associated with poor performance on heterophilic graphs [11, 64, 65], has been shown to be avoidable without altering the GNN architecture, simply by increasing the variance of weight initialization [66].
@@ 9.1 Interpretability
To create a parallelism with the interpretation of communities in PGMs and the output GNNs’ embeddings, we visualize the memberships and embeddings in a two-dimensional plan by applying a dimensionality reduction to these vectors using t-SNE [72]. We use a synthetic network as example, as we have access to ground truth (GT) communities. We first notice how nodes that have a more pronounced mixed-membership are placed on the border of their main community, closer to the other communities they also participate in, see Figure 5. This is somehow expected when plotting the GT $u_{i}$, given their shape. However, it is not trivial that the models are able to maintain this positional property in the inferred parameters. We find that both MULTITENSOR and GAT are able to place more mixed nodes closer to the border, as observed for the GT. For instance, nodes that belong to a large extent to two communities, are placed in between the two main clusters. However, while mixed-membership PGMs allow for a direct estimation of the mixed-membership, as this is their natural output, it is not clear how to extract this information automatically and in a principled way (beside visual inspection) in GNNs from their output embeddings, as they lack the innate ability to represent these memberships explicitly. To enhance the interpretability of GNNs, one promising direction is to develop methods that enable the extraction of interpretable mixed-membership information from their embeddings. Achieving this, however, would require architectural modifications to GNNs, as done in [18].
@@ 9.2 Computational Complexity
Table 2: Computational complexities for the models under analysis. NN the number of nodes, $E$ the total number of edges, $E^{(2)}$ the number of edges between nodes that are connected via a 2-hop path, $d_{\max}$ the maximum degree across all nodes, $K$ the number of communities, $Z$ the number of attribute categories, $F$ the input feature dimension, $F^{\prime}$ the hidden layer dimension, and $L$ the number of GNN layers.
@@
H2GCN $O(E\cdot d_{\max}+L\cdot((E+E^{(2)})\cdot F^{\prime})+N\cdot F\cdot F^{\prime})$
@@ 10 Conclusions
However, an important advantage of GNNs lies in their flexibility in processing input features. Unlike PGMs, GNNs do not rely on ad-hoc design choices to handle heterogeneous features, thus simplifying their implementation and widening their usage. Nevertheless, recent work [73] suggests that it may be possible to benefit from similar black-box tools as those used in GNNs to incorporate arbitrary types of attributes as input to PGMs.
Moreover, previous studies have shown that linear models can outperform GNNs in link prediction tasks [74, 75]. Given these limitations, GNNs should ideally be designed to learn automatically whether to discard irrelevant features or to select what features to use, potentially incorporating structural information such as the adjacency matrix of the graph. In the absence of this additional learning step, care should be taken when choosing what features to give in input.
@@ Appendix A Additional plots
In this section, we present two additional plots complementing the input features experiments presented in the main text: i) in Figure 7 we compare the performance of PGM and GNN models when using structure-based input features on both real and synthetic datasets; ii) in Figure 8 we further illustrate the performance of GNNs when fed either the full input feature or the clustered one on both real and synthetic datasets.
As shown before in Figure 1, from Figure 7 we can see that the performance of GNNs is generally comparable to that of PGMs when feeding structure-based input features.
@@
Furthermore, as shown previously in Figure 2, we observe a significant drop in GNN performance when the input features are clustered, and this effect is consistent across most datasets. The difference in performance between full and clustered features tends to be smaller when the input is already low-dimensional, as in the case of Peptides ($F=9$), or when the full feature performance on the dataset is already low, such as in the disassortative synthetic setting.
Figure 8: High vs. low dimensional features. AUC scores of link prediction in real and synthetic datasets when giving in input either the original feature of dimension $F$ (Full) or the one clustered with K-means (Clustered), i.e. $F=1$. For the PGMs, MTCOV takes one-dimensional (Clustered) features by design and the other two PGMs no feature at all. Bars and errors are mean and standard deviation over 5 cross-validation folds.
Appendix B Synthetic dataset details
We generate the synthetic networks following the generative modeling provided by MULTITENSOR [31]. The network consists of $N$ nodes and its connectivity is described by an adjacency matrix $A$, where $A_{ij}$ represents the number of edges from node $i$ to node $j$. The networks are generated probabilistically, assuming an underlying structure consisting of $K$ overlapping groups. Each node belongs to each group to an extent described by a $K$-dimensional membership vector. In order to generate directed networks, each node $i$ has two membership vectors, ${u}_{i}$ and ${v}_{i}$, which determine how $i$ forms outgoing and incoming links, respectively. When modeling undirected networks, we set ${u}_{i}={v}_{i}$. Then, the density of edges within and across communities is regulated by a $K\times K$-dimensional affinity matrix w{w}.
Given the parameters ${u}_{i},{v}_{j},{w}$, we sample entries $A_{ij}$ of the adjacency matrix from a Poisson distribution as:
$A_{ij}$ $\sim\text{Pois}(M_{ij})$ (3)
$M_{ij}$ $=\sum_{k,\ell=1}^{K}u_{ik}v_{j\ell}w_{k\ell}\,.$ (4)
In our current applications, we concentrate on the sparse case, where $M_{ij}$ is small, and we then clip $A_{ij}$ to consider only binary entries $A_{ij}\in\left\{0,1\right\}$.
To generate networks with either assortative or disassortative community structure, we control the form of the affinity matrix w{w}, which governs the interaction strengths between latent communities. In the assortative setting, w{w} is generated as a one-dimensional array of size $K$, corresponding to a diagonal affinity matrix. This implies that nodes are more likely to connect within their own community, resulting in higher intra-community edge density. In contrast, for disassortative networks, w{w} is sampled as a full $K\times K$ matrix, where the off-diagonal entries enable significant interactions between different communities. The entries of w{w} are independently drawn from a Gamma distribution with shape and rate parameters both equal to 1.0, i.e. $w_{kq}\sim\text{Gamma}(1.0,1.0)$. Node community memberships ${u}_{i}$ and ${v}_{j}$ are drawn from a symmetric Dirichlet distribution ${u}_{i},{v}_{j}\sim\text{Dirichlet}(\alpha\mathbf{1})$, with $\alpha=0.05$ ensuring that each node predominantly belongs to a small subset of communities; and $\mathbf{1}$ being a $K$-dimensional vector with entries equal to $1$.
To summarize, the parameters of the SBM are generated as:
${u}_{i},{v}_{i}$ $\sim\text{Dirichlet}(\alpha\mathbf{1})\ \text{with }\alpha=0.05,\quad\text{for each node }i=1,\ldots,N$ (5)
[…]
where $\delta_{k\ell}$ is the Kronecker delta ensuring a diagonal matrix in the assortative case and $a=b=1.0$ in both cases.
We generate directed networks of $N\in\left\{100,1000\right\}$ number of nodes; $K=5$ communities and average degree $\langle k\rangle\in\left\{5,20\right\}$. For each combination of size and average degree, we generate one homophilic and one heterophilic network. We sample $10$ networks for each parameter configuration.